Circular No. 8488

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005aj AND 2005ak Further to IAUC 8486, T. Puckett and J. Newton report the discovery

of an apparent supernova (mag 18) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Feb. 18.02 UT (and confirmed on frames taken by G. Sostero, Remanzacco, Italy, with a 0.25-m reflector on Feb. 18.87, the magnitude also reported as 18). The new object is located at α = 2^{h}58^{m}48^{s}.85, δ = +75^{o}45′13′′.8 (equinox 2000.0), which is 8′′.6 east and 51′′.7 north of the center of UGC 2411. Nothing is present at this location on images taken by Puckett on 2002 Feb. 11 and 2003 Feb. 21 (limiting mag ∼ 20.0).

Puckett adds that he and B. Kerns have discovered another apparent supernova (mag 15.7) on an unfiltered CCD frame taken with the 0.60m automated supernova patrol telescope on Feb. 23.41 UT. SN 2005ak is located at α = 14^{h}40^{m}32^{s}.06, δ = +3^{o}30′37′′.1 (equinox 2000.0), which is 1′′.9 east and 2′′.9 north of the center of the apparent host galaxy. Nothing is present at this location on images taken by Puckett on 2002 Feb. 13 and 2003 Apr. 1 (limiting mag ∼ 20.0). SN 2005ak was confirmed on CCD frames taken by A. Sehgal, Osoyoss, BC, with a 0.5-m reflector on Feb. 24.45 at mag 15.7.

SUPERNOVA 2005al IN NGC 5304 Further to IAUC 8482, C. Jacques reports the discovery of an apparent

supernova (mag 15.1) by E. Pimentel, T. Napoleao, C. Colesanti, and himself on unfiltered CCD images taken at Belo Horizonte, Brazil, on Feb. 24.16 and 26.95 UT. SN 2005al is located at α = 13^{h}50^{m}00^{s}.33, δ = −30^{o}34′34′′.2 (equinox 2000.0), which is 14′′.8 west and 7′′.5 north of the center of NGC 5304. Nothing is visible at this location on a CCD image taken on Feb. 15.29 (limiting mag 18.3) or on a Digitized Sky Survey image from 1975 (limiting red mag 20.0).

COMETS C/2005 B3 AND C/2005 B4 (SOHO) Continuation to IAUC 8487 (C/2005 B3 is a Kreutz sungrazer, while

C/2005 B4 belongs to the Meyer group):

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 B3 Jan. 26.821 20^{h}53.^{m}6 −20^{o}26′ C3 XL 2005-D18
C/2005 B4  30.688 20 56.6 −16 11 C2 BZ 2005-D18
2005 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green